OpenStudy (anonymous):

help please 5 over 21 divided by 15 over 7

OpenStudy (aaronandyson):

\[\frac{ 5 }{ 21 } \div \frac{ 15 }{ 7 }\]

OpenStudy (k_lynn):

\[\frac{ 5 }{ 21 }\div \frac{ 15 }{ 7 }\] That?

OpenStudy (anonymous):

yea just like that

OpenStudy (aaronandyson):

Reciprocate it I mean do this \[\frac{ 5 }{ 21 } \div \frac{ 7 }{ 15 }\]

OpenStudy (aaronandyson):

I think that div signs changes to multiply @k_lynn I am right or wrong?

OpenStudy (anonymous):

I thought when dividing fractions you cross multiply but I'm not so good with math lol so i might be wrong

OpenStudy (k_lynn):

To divide fractions, you can actually multiply by the reciprocal. So like this. \[\frac{ 5 }{ 21 }\times \frac{ 7 }{ 15 }\]Now all you need to do is multiply the numerators and then multiply the denominators. And yes, you're right @AaronAndyson :)

OpenStudy (aaronandyson):

So now What is 7 times 3? and 5 times 15?

OpenStudy (k_lynn):

You don't cross multiply, because you know all the numbers. If you had something that looked like this... \[\frac{ x }{ 34 }=\frac{ 2 }{ 7 }\]Then you would cross multiply.

OpenStudy (k_lynn):

So now to finish the question, all you need to do is this. \[\frac{ 5\times7 }{ 21\times15 }\]

OpenStudy (anonymous):

35/ 315 is the answer i got is that correct

OpenStudy (aaronandyson):

O.O Try again.

OpenStudy (aaronandyson):

7 times 3 is 21 5 times 3 is 15

OpenStudy (anonymous):

im confused lol

OpenStudy (aaronandyson):

\[5 \times 7 \over 21 \times 15\]

OpenStudy (aaronandyson):

is

OpenStudy (aaronandyson):

\[1 \times 1 \over 3 \times 3\]

OpenStudy (aaronandyson):

as 7 times 3 is 21 5 times 3 is 15

OpenStudy (aaronandyson):

so we just just do this \[3 \times 3\]

OpenStudy (aaronandyson):

what is 3 times 3?

OpenStudy (anonymous):

i did that and i got 35 for 5 * 7 and 315 for 21 *15 i don't get were i went wrong I'm confused sorry I'm really bad at math :/

OpenStudy (aaronandyson):

Okay 35/315 is ?

OpenStudy (aaronandyson):

35 times what is 315?

OpenStudy (aaronandyson):

You're method is correct but mine is a bit simple.

OpenStudy (aaronandyson):

35 times 9 is 315 Hence the answer is 9

OpenStudy (anonymous):

35 * 9 = 315

OpenStudy (aaronandyson):

Correct! So what is the answer?

OpenStudy (anonymous):

@AaronAndyson

OpenStudy (anonymous):

ohhh okay thank you

OpenStudy (aaronandyson):

We have 35/315 we know that 35 times 9 is 315 so can you tell me the answer now?

OpenStudy (anonymous):

so it would be 35 over 9 or just 9

OpenStudy (aaronandyson):

Just 9 as We divide both by 35.

OpenStudy (anonymous):

i got 1 over 9

OpenStudy (anonymous):

it was correct lol thank you

OpenStudy (aaronandyson):

No problem lol.